OpenStudy (anonymous):

need help with system: 3+2x-y=0 -3x-7y=10x

OpenStudy (anonymous):

Hellooooooooo?

OpenStudy (anonymous):

Solve for y in both equations, then, because y = h and y = j, h must be equal to j. Solve for x, then substitute x in for one of the original equations, and solve for y.

OpenStudy (anonymous):

y=2x+3 -3x-14x-21=10x

OpenStudy (anonymous):

27x=-21

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

solve it out.

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

isolate x out of 27x=21 then put that value into the x in y=2x+3

OpenStudy (anonymous):

if u dont understand how i got to that point, ask

OpenStudy (anonymous):

i see now...i think

OpenStudy (anonymous):

when in doubt ask.

OpenStudy (anonymous):

is it y=122/27?

OpenStudy (anonymous):

42/27+81/27=123/27

OpenStudy (anonymous):

3+2x-y=0 Move all terms not containing y to the right-hand side of the equation. -y=-2x-3 Multiply each term in the equation by -1. -y*-1=-2x*-1-3*-1 Multiply -y by -1 to get y. y=-2x*-1-3*-1 Simplify the right-hand side of the equation by multiplying out all the terms. y=2x+3

OpenStudy (anonymous):

-3x-7y=10x Since -3x does not contain the variable to solve for, move it to the right-hand side of the equation by adding 3x to both sides. -7y=3x+10x Simplify the right-hand side of the equation. -7y=13x Divide each term in the equation by -7. -(7y)/(-7)=(13x)/(-7) Simplify the left-hand side of the equation by canceling the common factors. y=(13x)/(-7) Simplify the right-hand side of the equation by simplifying each term. y=-(13x)/(7)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

ela its easier if you move y to the right from the getgo. left and right dont matter

OpenStudy (anonymous):

you are right, but for most simple people it's easier to stay with what they know. but for me it works either way.

OpenStudy (anonymous):

a=b b=a tomato tomato